# Hyperbolic recurrent neural network as the first type of non-Euclidean neural quantum state ansatz

### v2

@@ 6 Concluding remarks
In this work, we introduce the hyperbolic GRU as the first type of non-Euclidean neural quantum state (NQS) to be used in the variational Monte Carlo method (VMC) to approximate the ground state energy in quantum many-body physics, which are represented by the prototypical settings of the one-dimensional & two-dimensional transverse field Ising model (1D/2D TFIM) and the one-dimensional Heisenberg $J_{1}J_{2}$ $J_{1}J_{2}J_{3}$ models. Our results show an interesting trend regarding the performance of the hyperbolic GRU as benchmarked against the performances of its Euclidean counterpart. In particular, using either the direct training or inference results, we note the following.
@@
  • In the 1D TFIM setting, with the number of spins $N$ being 20, 40, 80 and 100, where the Hamiltonian comprises only the nearest neighbor interaction and a transverse magnetic field, with no hierarchical interaction structure, three out of four times, hyperbolic GRU does not outperform either Euclidean RNN or GRU.
@@
  • In the 2D TFIM setting, with the number of size of the 2D square lattice $(N_{x},N_{y})$ being $(5,5),(7,7),(8,8)$, $(9,9)$ corresponding to $N=25,49,64,81$ spins, where the Hamitonian comprises the horizontal and vertical nearest neighbor interaction in two dimensions and a transverse magnetic field, 2D Euclidean RNN always emerge as the best ansatz, followed by 1D hyperbolic GRU, which almost always outperforms 1D Euclidean GRU, except for a single instance in the training result for $(N_{x},N_{y})=(8,8)$. In this case, due to the artificial rearrangement or reshaping/folding of the 1D spin chain to mimic the 2D lattice, a hierarchy in the interaction structure, comprising the first and the $N^{th}$ nearest neighbor interactions, is introduced when originally faraway spins (at site $i$ and site $i+N$) in the 1D chains become immediate vertical neighbors. The fact that the hyperbolic GRU NQS ansatz outperformed the Euclidean GRU NQS ansatz for this setting is our first clue of the role played by the hyperbolic geometry of the neural network in determining its performance when there is a hierarchical structure in the Hamiltonian.
@@
  • In the 1D Heisenberg $J_{1}J_{2}$ setting with $N=50$ spins, $J_{1}=1.0$ (fixed) and $J_{2}=0.0,0.2,0.5,0.8$, hyperbolic GRU definitively outperforms Euclidean GRU three out of four times when the $J_{1}J_{2}$ Hamiltonian exhibits a hierarchical interaction structure when $J_{2}\neq 0$ which takes into account the second nearest neighbor interactions in addition to the first nearest neighbor interactions. This is our second clue of the role played by the hyperbolic geometry of the neural network.
@@
  • In the 1D Heisenberg $J_{1}J_{2}J_{3}$ setting with $N=30$ spins, $J_{1}=1.0$ (fixed) and $(J_{2},J_{3})=(0.0,0.5),(0.2,0.2)$, $(0.2,0.5)$, $(0.5,0.2)$, hyperbolic GRU definitively outperforms Euclidean GRU four out of four times in the first set of experiments (where the networks have the same number of parameters), and three out of four times in the second set of VMC experiments (where the hyperbolic GRU network has around 10-20 percent fewer parameters than the Euclidean GRU network). In this setting, the $J_{1}J_{2}J_{3}$ Hamiltonian exhibits even more pronounced hierarchical interaction structures comprising the first, second and third nearest neighbors. This is our third clue.
@@
Given these findings, and the established results in natural language processing that hyperbolic RNNs almost always outperform their Euclidean versions when the training data exhibit hierarchical, tree-like structures [30], [35], [33], it is our hypothesis that hyperbolic RNN-based NQS will likely outperform Euclidean RNN-based NQS in all quantum spin systems in which there is a hierarchical structure in the interaction part of the Hamiltonian. More works are needed to confirm or refute this hypothesis. If true, this would be an exciting new direction opened up by the possibility of better performances delivered hyperbolic and potentially other non-Euclidean NQS in the problem of approximating the ground state energy of quantum many-body Hamiltonians using VMC - a problem that has been studied using exclusively conventional or Euclidean neural networks.
@@
Appendix A Convergence curves
A.1 1D TFIM
Figure 11: Convergence curves of the mean energy from the VMC experiments for 1D TFIM with various $N$ (clockwise from top: $N=20,40,100,80$ spins) with the three types of NQS ansatzes being eRNN-50-s50, eGRU-50-s50, hypGRU-50-s50. Each subfigure contains a zoomed-in plot showing a narrower range of energy values close to the exact DMRG value.
Figure 12: Convergence curves of the energy variances using eRNN-50-s50, eGRU-50-s50, hypGRU-50-s50 in VMC runs for 1D TFIM with various $N$ (clockwise from top: $N=20,40,100,80$ spins). The red dashed line corresponds to the line of zero variance. Each subfigure contains a zoomed-in plot showing a narrower range of variance values close to the zero line.
A.2 2D TFIM
@@ Appendix A Convergence curves
Figure 14: Convergence curves of the energy variances using 2D Euclidean RNN and 1D Euclidean/hyperbolic GRU ansatzes for the 2D TFIM with ($N_{x},N_{y}$) = (5,5), (7,7), (9,9), (8,8) lattices (clockwise from top). The red dashed line corresponds to the line of zero variance. Each subfigure contains a zoomed-in plot showing a narrower range of variance values close to the zero line.
A.3 1D Heisenberg $J_{1}J_{2}$ model
Figure 15: Convergence curves of the mean energy using Euclidean GRU and hyperbolic GRU from VMC runs of 1D Heisenberg $J_{1}J_{2}$ model with different $J_{2}$ values - clockwise from top, $J_{2}$ = 0.0, 0.2, 0.8, 0.5 - for $N=50$ spins.
Figure 16: Convergence curves of the energy variances using Euclidean GRU and hyperbolic GRU from VMC runs of 1D Heisenberg $J_{1}J_{2}$ model with different $J_{2}$ values - clockwise from top, $J_{2}$ = 0.0, 0.2, 0.8, 0.5 - for $N=50$ spins. The red dashed line corresponds to the line of zero variance.
A.4 1D Heisenberg $J_{1}J_{2}J_{3}$ model
@@
Figure 18: Convergence curves of the energy variances using Euclidean GRU and hyperbolic GRU from VMC runs of 1D $J_{1}J_{2}J_{3}$ Heisenberg model with different $(J_{2},J_{3})$ values - clockwise from top, ($J_{2},J_{3}$) = (0.0,0.5), (0.2,0.2), (0.5,0.2), (0.2,0.5)- for $N=30$ spins. The red dashed line corresponds to the line of zero variance.
@@
Figure 20: Convergence curves of the energy variances using Euclidean GRU and hyperbolic GRU from VMC runs of 1D $J_{1}J_{2}J_{3}$ Heisenberg model with different $(J_{2},J_{3})$ values - from top to bottom, ($J_{2},J_{3}$) = (0.2,0.5), (0.5,0.2) - for $N=30$ spins. The red dashed line corresponds to the line of zero variance. Each subfigure contains a zoomed-in plot showing a narrower range of variance values close to the zero line.

### v3

@@ 6 Concluding remarks
In this work, we introduce the hyperbolic GRU as the first type of non-Euclidean neural quantum state (NQS) to be used in the variational Monte Carlo method (VMC) to approximate the ground state energy in quantum many-body physics, which are represented by the prototypical settings of the one-dimensional & two-dimensional transverse field Ising model (1D/2D TFIM) and the one-dimensional Heisenberg $J_{1}J_{2}$ & $J_{1}J_{2}J_{3}$ models. Our results show an interesting trend regarding the performance of the hyperbolic GRU as benchmarked against the performances of its Euclidean counterpart. In particular, from the obtained results of the experiments performed in this work, we note the following.
@@
  • In the 1D TFIM setting, with the number of spins $N$ being 20, 40, 80 and 100, where the Hamiltonian comprises only the nearest neighbor interaction and a transverse magnetic field, with no hierarchical interaction structure, hyperbolic GRU showed a comparable performance to Euclidean GRU, with both GRU variants outperforming the Euclidean RNN variant.
@@
  • In the 2D TFIM setting, with the number of size of the 2D square lattice $(N_{x},N_{y})$ being $(5,5),(7,7),(8,8)$, $(9,9)$ corresponding to $N=25,49,64,81$ spins, where the Hamitonian comprises the horizontal and vertical nearest neighbor interaction in two dimensions and a transverse magnetic field, 1D hyperbolic GRU always outperformed 1D Euclidean GRU (while the best performing ansatz, the 2D Euclidean RNN NQS, was included purely to provide a benchmark). In this case, due to the artificial rearrangement or reshaping/folding of the 1D spin chain to mimic the 2D lattice, a hierarchy in the interaction structure, comprising the first and the $N^{th}$ nearest neighbor interactions, is introduced when originally faraway spins (at site $i$ and site $i+N$) in the 1D chains become immediate vertical neighbors. The fact that the hyperbolic GRU NQS ansatz outperformed the Euclidean GRU NQS ansatz for this setting is our first clue of the role played by the hyperbolic geometry of the neural network in determining its performance when there is a hierarchical structure in the Hamiltonian.
@@
  • In the 1D Heisenberg $J_{1}J_{2}$ setting with $N=50$ spins, $J_{1}=1.0$ (fixed) and $J_{2}=0.0,0.2,0.5,0.8$, hyperbolic GRU definitively outperformed Euclidean GRU when the $J_{1}J_{2}$ Hamiltonian exhibits a hierarchical interaction structure - when $J_{2}\neq 0$, which takes into account the second nearest neighbor interactions in addition to the first nearest neighbor interactions. This is our second clue of the role played by the hyperbolic geometry of the neural network.
@@
  • In the 1D Heisenberg $J_{1}J_{2}J_{3}$ setting with $N=30$ spins, $J_{1}=1.0$ (fixed) and $(J_{2},J_{3})=(0.0,0.5),(0.2,0.2)$, $(0.2,0.5)$, $(0.5,0.2)$, hyperbolic GRU again definitively outperformed Euclidean GRU in four out of four cases in the first set of experiments (where the networks have the same number of parameters). In the second set of experiments (involving only the two cases $J_{2}=(0.2,0.5)$ and $J_{2}=(0.5,0.2)$), where the hyperbolic GRU network has around 10-20 percent fewer parameters than the Euclidean GRU network, hyperbolic GRU attained a comparable performance with Euclidean GRU in one case and outperformed Euclidean GRU in another case. In this setting, the $J_{1}J_{2}J_{3}$ Hamiltonian exhibits even more pronounced hierarchical interaction structures comprising the first, second and third nearest neighbors. This is our third clue.
@@
Given these findings, and the established results in natural language processing that hyperbolic RNNs almost always outperform their Euclidean versions when the training data exhibit hierarchical, tree-like structures [30], [35], [33], it is our hypothesis that hyperbolic RNN-based NQS will likely outperform Euclidean RNN-based NQS in all quantum spin systems in which there is a hierarchical structure in the interaction part of the Hamiltonian. Given the fact that the experiments done in this work have only involved small-size systems due to our limited computing resources, it is obvious that more works are needed to either confirm or refute this hypothesis. If true, this would be an exciting new direction opened up by the possibility of better performances delivered hyperbolic and potentially other non-Euclidean NQS in the problem of approximating the ground state energy of quantum many-body Hamiltonians using VMC - a problem that has been studied using exclusively conventional or Euclidean neural networks.
@@
Appendix A Two-dimensional RNN model
where $f$ is a nonlinear activation function, the input $\vec{x}_{t}=(\vec{x}^{0}_{t},\vec{x}^{1}_{t})$ and hidden state $\vec{h}_{t}=(\vec{h}^{0}_{t},\vec{h}^{1}_{t})$ at timestep $t$ are now a 2D vectors comprising the horizontal and vertical components. Correspondingly, there is a doubling of the ususal weight matrices $U$ and $W$ to $U_{h},U_{v}$, $W_{h},W_{v}$. For a 2D spin configuration $\mathbf{{\sigma}}$ of dimension $(N_{x},N_{y}$), an entry ${\sigma}_{i,j}$ is autoregressively generated based on the two spins and two hidden states from its previously generated horizontal and vertical neighbors. For the 1D Euclidean/hyperbolic GRU, during the sampling process, the site ${\sigma}_{i,j}$ is autoregressively generated based on only the previously generated site in a raster manner shown in Fig.9. At each time step $t$ during the training phase, the 2D Euclidean RNN NQS ansatz generates 50 two-dimensional samples, each of which is a (2D) spin configuration (samples) of size $(N_{x},N_{y}$), while the 1D Euclidean/hyperbolic GRU ansatz generates 50 one-dimensional samples, each of which is a (1D) spin configuration of length $N_{x}\times N_{y}$. The computation of the energy for the 1D ansatz requires a reshaping of the 1D samples of length $N_{x}\times N_{y}$ into 2D samples of size $(N_{x},N_{y})$.
 1\,1\, 2\,2\, 3\,3\, 4\,4\, 5\,5\, 6\,6\, 7\,7\, 8\,8\, 9\,9\,1010111112121313141415151616
 1\,1\, 2\,2\, 3\,3\, 4\,4\, 8\,8\, 7\,7\, 6\,6\, 5\,5\, 9\,9\,1010111112121313141415151616
 1\,1\, $2\,\,\ldots\,$$1\,\,\ldots\,N+N\,N+N\,\,\ldots\,$(N1)N+N(N-1)N+N
[…]
Appendix B Convergence curves
B.1 1D TFIM
Figure 11: Convergence curves of the mean energy from the VMC experiments for 1D TFIM with various NN (clockwise from top: $N=20,40,100,80$ spins) with the three types of NQS ansatzes being eRNN-50-s50, eGRU-50-s50, hypGRU-50-s50. Each subfigure contains a zoomed-in plot showing a narrower range of energy values close to the exact DMRG value.
B.2 2D TFIM
@@ Appendix B Convergence curves
B.3 1D Heisenberg $J_{1}J_{2}$ model
B.3.1 Results without gradient clipping
In this section, we recorded the results obtained from training the models without using gradient clipping.
Ansatz $J_{2}=0.0$ $J_{2}=0.2$ $J_{2}=0.5$ $J_{2}=0.8$
Euclidean GRU -21.2469 -17.6566 -17.3347 -19.3759
(0.0094) (0.0188) (0.0128) (0.0133)
Hyperbolic GRU -20.4145 -19.7781 -18.6643 -20.0535
(0.0112) (0.0113) (0.0066) (0.0120)
Exact (DMRG) -21.9721 -20.3150 -18.7500 -20.9842
Table 11: VMC inference results (using $10^{4}$ samples) for the 1D $J_{1}J_{2}$ system with the Hamiltonian Eq.(53) with $N=50$, $J_{1}=1$ and $J_{2}=0.0,0.2,0.5,0.8$. For each ansatz, the mean $E$ is listed first, together with the standard error (in brackets). The best result at each $J_{2}$ value is noted in bold. These results were obtained by training the NQS without gradient clipping.
Figure 13: Comparisons of the performances of Euclidean and hyperbolic GRU NQS ansatzes (trained without gradient clipping) listed in Table 5 from VMC runs of 1D $J_{1}J_{2}$ model with different $J_{2}$ values (clockwise from top: $J_{2}$ =0.0, 0.2, 0.8, 0.5) for $N=50$ spins. In each subfigure, the mean energy for each NQS ansatz is shown as a dot with an error bar (representing the standard error). The specific variants of Euclidean/hyperbolic GRU used in each of the four cases are listed in Table 5. In this plot, we do not specify the exact variants, but label them simply as Euclidean GRU or hyperbolic GRU to emphasize the geometry of the ansatzes.
B.3.2 Convergence curves
Figure 14: Convergence curves of the mean energy using Euclidean GRU and hyperbolic GRU from VMC runs of 1D Heisenberg $J_{1}J_{2}$ model with different $J_{2}$ values. These curves were obtained from training with no gradient clipping. Clockwise from top, $J_{2}$ = 0.0, 0.2, 0.8, 0.5 - for $N=50$ spins.
Figure 15: Convergence curves of the mean energy using Euclidean GRU and hyperbolic GRU from VMC runs of 1D Heisenberg $J_{1}J_{2}$ model with different $J_{2}$ values, with gradient clipping during the training process, - clockwise from top, $J_{2}$ = 0.0, 0.2, 0.8, 0.5 - for $N=50$ spins.
Figure 16: Convergence curves of the mean energy of Euclidean GRU only from VMC runs of 1D Heisenberg $J_{1}J_{2}$ model with different $J_{2}$ values with and without gradient clipping during the training process. Curves obtained from training with no gradient clipping are marked by ‘-o’ in their names, while those obtained from training with gradient clipping are not. Clockwise from top, $J_{2}$ = 0.0, 0.2, 0.8, 0.5 - for $N=50$ spins.
Figure 17: Convergence curves of the mean energy of hyperbolic GRU only from VMC runs of 1D Heisenberg $J_{1}J_{2}$ model with different $J_{2}$ values with and without gradient clipping during the training process. Curves obtained from training with no gradient clipping are marked by ‘-o’ in their names, while those obtained from training with gradient clipping are not. For the case of $J_{2}=0.5$, it was not necessary to perform the experiment again using gradient clipping due to no instability present in the training. Clockwise from top, $J_{2}$ = 0.0, 0.2, 0.8, 0.5 - for $N=50$ spins.
B.4 1D Heisenberg $J_{1}J_{2}J_{3}$ model
@@ B.3.2 Convergence curves
Appendix C Data Availability Statement
The data and results (in the form of Python scripts, Jupyter notebooks, and trained Euclidean/hyperbolic neural network models) supporting the findings of this work are available at the following Github repository: https://github.com/lorrespz/nqs_hyperbolic_rnn